# DecompDreamer: Advancing Structured 3D Asset Generation with Multi-Object Decomposition and Gaussian Splatting

### v1

@@ 5.2 Ablation Study
Training Routine. We conduct an ablation study to assess the impact of different components in our training routine. Figure 6 compares several variations: ”Base” directly trains on the input prompt without decomposition using SD3.5, failing to generate a coherent 3D structure. $\mathcal{L}_{joint}$ optimizes only relationships, forming a well-structured scene but with poor disentanglement, causing objects to blend. $\mathcal{L}_{joint}+\sum\mathcal{L}_{obj}$ assigns equal weight to relationships and objects, however it confuses the model, leading to structural inconsistencies. For instance, a knight riding a horse appears structurally different from a standalone knight, confusing the model and resulting in poor relationship modeling. In contrast, our progressive optimization strategy, which prioritizes relationships before refining individual objects, significantly enhances structural integrity. The first stage ensures well-formed relationships, while the second stage improves object quality while maintaining these relationships. Additional results and analysis are in Section A.4 of the appendix.
Azimuth Adjustment. We conduct an ablation study to evaluate the impact of azimuth adjustment and object-view-dependent text prompts in (5). Without azimuth correction, objects exhibit noticeable disorientation. As shown in Figure 9, omitting object-view adjustments in the robot example results in misplaced features, such as eyes appearing on the back of its head. This misalignment occurs because the global front view differs from the object’s actual orientation, causing $\mathcal{L}_{obj}$ and $\mathcal{L}_{joint}$ to optimize in conflicting directions. Further results and analysis are in Section A.4 of the appendix.
Comparison with SD2.1-based DecompDreamer. To evaluate DecompDreamer’s optimization routine, we replace the FMD-based Stable Diffusion 3.5 with SDS-based Stable Diffusion 2.1. Compared to other SOTA methods, DecompDreamer with SD2.1 achieves better inter-object relationship modeling and successfully generates all objects in a scene. As shown in Figure 13, under the “hero dragon” prompt, methods like GraphDreamer and MVDream fail to generate all objects, while GALA3D struggles with relationship modeling (e.g., the dragon faces the wrong direction). In contrast, DecompDreamer with SD2.1 correctly captures both objects and their relationships. Further improvements are achieved by replacing SDS with Flow Matching Distillation (FMD), enhancing sharpness and reducing iteration requirements by 75%. Additional discussion are in Section A.4 of the appendix.
6 Conclusion
In this paper, we introduce DecompDreamer, a Gaussian splatting-based text-to-3D generation framework designed to handle compositional prompts with a large number of objects and intricate inter-object relationships. Unlike existing methods, DecompDreamer not only generates high-quality compositional 3D structures but also ensures coherent individual object generation, providing greater flexibility and control for downstream applications. Extensive experiments on prompts containing 2–11 objects validate DecompDreamer’s effectiveness in capturing complex relationships while maintaining object-level disentanglement. While our primary focus has been on modeling inter-object relationships, the generation quality of individual objects is still constrained by SDS optimization, making it susceptible to artifacts such as the Janus problem. However, recent advancements such as MT3D [27] and DreamControl [16] can be seamlessly integrated into our approach to enhance object fidelity if needed. Additionally, DecompDreamer is tailored for inter-object relationships and may not generalize well to decomposing a single entity into finer-grained components. Addressing these challenges may require a more structured, hierarchical representation for single-object decomposition, which we leave for future exploration.
@@ Appendix A Appendix
A.1 Notations and Algorithm
@@
$N$ Number of objects in the scene
@@
$\mathcal{L}_{\text{obj}}$ Object-specific FMD loss
@@
$\mathcal{L}_{\text{FMD}}$ Flow Matching Distillation loss
@@
Algorithm 1 Structured Decomposition and Progressive FMD Optimization
@@
2:$\mathcal{I},G,N\leftarrow$ VLM($t^{g}$) $\triangleright$ Generate image and scene graph
3:$E,O,\{t^{o}_{i},t^{e}_{ij},x_{i},y_{i},z_{i},s_{i},\theta_{i}\}\leftarrow$ VLM($\mathcal{I},G$)
4:for $i=1$ to $N$ do $\triangleright$ Convert objects to Gaussians
@@
9:$\mathcal{G}=\bigcup_{i=1}^{N}\mathcal{G}_{o_{i}}$triangleright Unified Gaussian representation
10:Stage 2: Progressive FMD Optimization
11:$n_{e}\leftarrow|E|$triangleright Number of edges
12:for $t=1$ to $T_{\text{max}}$ do
13:     if $t<\gamma.T_{\text{max}}$ then $\triangleright$ Joint Optimization
14:         if $t\bmod(n_{e}+1)\neq 0$ then
15:              $i,j\leftarrow E[t\bmod n_{e}]$
16:              $\mathcal{G}^{\prime}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i},\quad\mathcal{G}^{\prime}_{o_{j}}\leftarrow\mathcal{G}_{o_{j}}+T^{o}_{j}$
17:              $\mathcal{L}_{\text{joint}}\leftarrow\nabla_{\mathcal{G}^{\prime}_{o_{i}},\mathcal{G}^{\prime}_{o_{j}}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij})$
@@
18:              $\mathcal{L}_{\text{obj}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{FMD}}(z;y^{o}_{a_{O}})$
19:              $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{joint}}+\lambda\left(\frac{t}{T_{\text{max}}}\right)^{2}\mathcal{L}_{\text{obj}}$
20:         else
21:              for $i=1$ to $N$ do
22:                  $\mathcal{G}^{\prime}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i}$
23:              end for
24:              $\mathcal{L}_{\text{scene}}\leftarrow\nabla_{\mathcal{G}}\mathcal{L}_{\text{FMD}}(z;y^{g})$
25:              $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{scene}}$
26:         end if
27:     $\triangleright$ Targeted Optimization
28:         $i,j\leftarrow E[t\bmod n_{e}]$
29:         $\mathcal{L}_{\text{target}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij})$
30:         $\mathcal{L}_{\text{obj}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{FMD}}(z;y^{o}_{a_{O}})$
31:         $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{target}}+\mathcal{L}_{\text{obj}}$
32:     end if
@@
33:     $\mathcal{G}\leftarrow$ Adam($\nabla\mathcal{L}_{\text{loss}}$) $\triangleright$ Update step
34:     if IsRefinementIteration($t$then
35:         $\mathcal{G}\leftarrow$ DensifyAndPrune($\mathcal{G}$)
36:     end if
37:end for
A.2 Hyperparameters
Our optimization framework follows the learning rates and camera parameter updates used in LucidDreamer [20]. Each object is allocated 1500 iterations, with the first 450 iterations serving as a warm-up. Densification and pruning begin after 100 iterations and continue until 900 iterations, with a 100-iteration interval for both operations. The azimuth angle is constrained to [-180, 180] to ensure stable optimization. The minimum timestep is 0.02, while the maximum timestep, initially 0.98, is annealed to 0.5 during warm-up and remains constant thereafter. The hyperparameters $\lambda$ and $\gamma$ in Eq. 6 are set to $8$ and $0.6$, respectively. The warm-up phase is restarted in the second stage for 30% of the remaining iterations.
A.3 Coarse Initialization Details
Each object $O_{i}$ is initialized with a point cloud $P(O_{i})$ generated using Point-E [28]. To ensure proper alignment within the scene, each point cloud undergoes a transformation based on spatial information predicted by the VLM. Let $(x_{i},y_{i},z_{i})$ denote the center coordinates, $s_{i}$ the relative size, and $\theta_{i}$ the orientation of object $O_{i}$, then the alignment transformation is given by:
$P_{\text{scene}}=\bigcup_{i=1}^{n}P(O_{i})_{\text{aligned}}$
$\mathcal{G}=\bigcup_{i=1}^{n}\mathcal{G}_{o_{i}}$
A.4 Additional Ablation Studies
@@
Azimuth Adjustment We further analyze the effect of azimuth adjustment and object-view-dependent text prompts in multi-object scenes. Without azimuth correction, global view-dependent prompts can lead to conflicting optimizations, as different objects may have different front-facing orientations. In Figure 9, incorrect azimuth alignment results in misplaced features, highlighting the issue of optimizing objects using a fixed global view. By applying object-specific azimuth corrections, we ensure that each object is optimized relative to its correct orientation, leading to better spatial alignment and disentanglement.
@@
Comparison with Stable Diffusion 2.1 We further analyze the impact of using SDS-based SD2.1 versus FMD-based SD3.5 in DecompDreamer. While SDS-based SD2.1 outperforms other SOTA methods in relationship modeling, it still struggles with sharpness and convergence speed. By replacing SDS with FMD, we significantly improve object detail and reduce optimization time, requiring only one-fourth of the iterations. As shown in Figure 13, FMD produces sharper edges and more refined object details compared to SDS.
@@
A.5 Sample usage of ChatGPT
A.6 More Comparisons
Figure 15: More qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D [43], GraphDreamer [13], MVDream [36], LucidDreamer [20], Magic3D [21] and HiFA [44].
Figure 16: More qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D [43], GraphDreamer [13], MVDream [36], LucidDreamer [20], Magic3D [21] and HiFA [44].

### v2

@@ 6.3 Ablation Study
We conduct an ablation study to validate our design choices, with results in Figure 7. The Base model, trained without decomposition, fails to produce coherent 3D structures (e.g., a malformed horse). $\mathcal{L}_{\text{joint}}$ optimizes inter-object relationships, resulting in better layout but poor disentanglement. For instance, the 3D object corresponding to the horse includes a leg from the knight. $\mathcal{L}_{\text{joint}}+\sum\mathcal{L}_{\text{obj}}$ applies equal weight to both relational and object-level optimization. However, this confuses the model, often leading to structural inconsistencies where neither relationships nor individual objects are properly generated. For example, in the full composition, the horse contains only a faint leg from the knight, while the individual knight component lacks well-defined limbs. In contrast, our full composition-aware optimization succeeds. The staged curriculum first ensures a coherent relational structure before refining object fidelity, which resolves the conflicts observed in other variants and enhances overall structural integrity. Additional results are in Appendix A.7.
7 Conclusion and Future Works
In this paper, we reframed compositional text-to-3D generation as a problem of optimization scheduling, arguing that the failures of prior methods are predictable outcomes of heuristics that collapse under a combinatorial explosion of conflicting gradients. We introduced DecompDreamer, which addresses this theoretical bottleneck with a staged, composition-aware optimization strategy. Functioning as an implicit curriculum, our method temporally decouples competing objectives by first establishing a coherent relational scaffold before refining the high-fidelity details of individual objects. Our extensive experiments validate this approach, demonstrating state-of-the-art performance in fidelity, disentanglement, and spatial coherence, especially on complex scenes where other heuristics fail. While this work focuses on generating assets composed of distinct object types, a promising future direction is the decomposition of a single entity into finer-grained subcomponents—such as segmenting a 3D human model into regions like the face, hands, or limbs. In such settings, DecompDreamer’s composition-aware optimization could serve as a foundational strategy. Addressing these challenges may require improved representations for fine-grained parts, which we leave as an exciting avenue for future work.
@@ Appendix A Appendix
A.1 Hyperparameters
Our optimization framework follows the learning rates and camera parameter updates used in LucidDreamer Liang et al. (2024). Each object is allocated 1500 iterations, with the first 450 iterations serving as a warm-up. Densification and pruning begin after 100 iterations and continue until 900 iterations, with a 100-iteration interval for both operations. The azimuth angle is constrained to [-180, 180] to ensure stable optimization. The minimum timestep is 0.02, while the maximum timestep, initially 0.98, is annealed to 0.5 during warm-up and remains constant thereafter. The hyperparameters $\lambda$ and $\gamma$ in Eq. 6 are set to $8$ and $0.6$, respectively. The warm-up phase is restarted in the second stage for 30% of the remaining iterations.
A.2 Coarse Initialization Details
Each object $O_{i}$ is initialized with a point cloud $P(O_{i})$ generated using Point-E Nichol et al. (2022). To ensure proper alignment within the scene, each point cloud undergoes a transformation based on spatial information predicted by the VLM. Let $(x_{i},y_{i},z_{i})$ denote the center coordinates, $s_{i}$ the relative size, and $\theta_{i}$ the orientation of object $O_{i}$, then the alignment transformation is given by:
$P_{\text{scene}}=\bigcup_{i=1}^{n}P(O_{i})_{\text{aligned}}$
$\mathcal{G}=\bigcup_{i=1}^{n}\mathcal{G}_{o_{i}}$
A.3 Notations and Algorithm
@@
NN Number of objects in the scene
@@
$\mathcal{L}_{\text{obj}}$ Object-specific SDS loss
@@
$\mathcal{L}_{\text{SDS}}$ Score Distillation Sampling loss
@@
Algorithm 1 Structured Decomposition and Composition-Aware Optimization
@@
[…]
4:for $i=1$ to NN do $\triangleright$ Convert objects to Gaussians
@@
[…]
11:for $t=1$ to $T_{\text{translation}}$ do
@@
12:  $\mathcal{L}_{\text{spatial-error-correction}}\leftarrow\sum\limits_{(i,j)\in E}\nabla_{T^{O}_{i},T^{O}_{j}}\mathcal{L}_{\text{SDS}}(z;y^{e}_{ij})$
@@
13:  $T^{O}\leftarrow$ Adam($\nabla\mathcal{L}_{\text{spatial-error-correction}}$) $\triangleright$ Update step
14:end for
15:Stage 2: Composition-Aware Optimization
[…]
17:for $t=1$ to $T_{\text{max}}$ do
18:  if $t<\gamma.T_{\text{max}}$ then $\triangleright$ Joint Optimization
19:   if $t\bmod(n_{e}+1)\neq 0$ then
20:     $i,j\leftarrow E[t\bmod n_{e}]$
21:     $\mathcal{G}^{\prime}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i},\quad\mathcal{G}^{\prime}_{o_{j}}\leftarrow\mathcal{G}_{o_{j}}+T^{o}_{j}$
22:     $\mathcal{L}_{\text{joint}}\leftarrow\nabla_{\mathcal{G}^{\prime}_{o_{i}},\mathcal{G}^{\prime}_{o_{j}}}\mathcal{L}_{\text{SDS}}(z;y^{e}_{ij})$
23:     $\mathcal{L}_{\text{obj}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{SDS}}(z;y^{o}_{a_{O}})$
24:     $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{joint}}+\lambda\left(\frac{t}{T_{\text{max}}}\right)^{2}\mathcal{L}_{\text{obj}}$
25:   else
26:     for $i=1$ to NN do
27:      $\mathcal{G}^{\prime}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i}$
28:     end for
29:     $\mathcal{L}_{\text{scene}}\leftarrow\nabla_{\mathcal{G}}\mathcal{L}_{\text{SDS}}(z;y^{g})$
30:     $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{scene}}$
31:   end if
[…]
33:   $i,j\leftarrow E[t\bmod n_{e}]$
34:   $\mathcal{L}_{\text{target}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{SDS}}(z;y^{e}_{ij})$
35:   $\mathcal{L}_{\text{obj}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{SDS}}(z;y^{o}_{a_{O}})$
36:   $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{target}}+\mathcal{L}_{\text{obj}}$
37:  end if
38:  $\mathcal{G}\leftarrow$ Adam($\nabla\mathcal{L}_{\text{loss}}$) $\triangleright$ Update step
39:  if IsRefinementIteration($t$) then
40:   $\mathcal{G}\leftarrow$ DensifyAndPrune($\mathcal{G}$)
41:  end if
42:end for
A.4 Execution time comparison between optimization-based methods
Table 3 presents the time (in minutes) required by each method to generate 3D content as the number of objects increases. Our method demonstrates a consistent and scalable performance trend—requiring approximately 3/4 the time of GALA3D across object counts, while producing higher-fidelity, geometrically consistent, and semantically accurate compositions. In contrast, GraphDreamer maintains a flat generation time of 180 minutes but fails to scale beyond 4 objects, often struggling to handle increased relational complexity.
Methods such as LucidDreamer, Magic3D, MVDream, and Trellis exhibit nearly constant and fast generation times (ranging from 0.2 to 49 minutes); however, they are primarily designed for single-object synthesis. While they can sometimes handle two-object scenes and occasionally succeed with three, they consistently fail to capture complex inter-object relationships and compositional semantics when the number of entities exceeds this threshold. Our method is the only approach that scales gracefully in both visual fidelity and semantic alignment, while remaining significantly faster than GALA3D and more robust than GraphDreamer at higher object counts.
Table 3: Time (in minutes) comparison of different methods across varying number of objects.
Method #2 #3 #4 #6 #11 Magic3D 49 49 49 49 49 GraphDreamer 180 180 180 - - GALA3D 120 180 240 360 660 LucidDreamer 45 45 45 45 45 MVDream 43 43 43 43 43 Trellis 0.2 0.2 0.2 0.2 0.2 Ours 90 135 180 270 495
A.5 User Study Protocol
Our user study was conducted under a strict double-blind and randomized protocol to ensure fair and unbiased comparisons. For each of the 30 prompts, 20 human evaluators were presented with the set of rendered videos generated by our method alongside all baseline methods. Participants were given the following clear instruction: "Please watch each set of videos and select the one you think best matches the given prompt. Please base your selection on both image quality and semantic alignment." To prevent any ordering bias, the sequence of videos shown for each prompt was randomized for every evaluator. Furthermore, the methods used to generate the videos were anonymized, meaning participants were unaware of which result corresponded to which method. This double-blind setup ensures that selections were based solely on the perceived quality and faithfulness of the result.
A.6 Quantitative Comparison and Discussion
To evaluate the performance of DecompDreamer, we employ four complementary metrics: CLIP Score Radford et al. (2021), Pick-a-Pic Kirstain et al. (2023), Text-to-3D Alignment Duggal et al. (2025), and a user study, applied across 30 prompts containing 2–11 objects. CLIP Score measures the semantic similarity between rendered images and input prompts using vision-language embeddings, providing a proxy for text-image alignment. Pick-a-Pic is a learned perceptual metric that models human preferences between image pairs conditioned on a prompt, capturing fidelity and relevance. Text-to-3D Alignment quantitatively assesses object presence, count, and spatial relations using grounding-based techniques over rendered views. The user study collects human judgments on object identity, relational correctness, and overall scene plausibility. As shown in Table 1, DecompDreamer achieves the highest scores across all four metrics. It leads in CLIP Score (34.5%) and Pick-a-Pic accuracy (24.0%), with strong generalization to complex scenes. It also attains the best Text-to-3D Alignment score (73.4%) and is preferred by users in 61.0% of cases overall, reaching 73.0% in high-object-count settings. These results highlight DecompDreamer’s ability to generate semantically faithful, perceptually preferred, and structurally coherent 3D content.
While DecompDreamer achieves the highest scores across all automated metrics, the observed performance gap—typically around 5%—is narrower than expected. This is largely attributable to the limitations of current evaluation metrics in capturing fine-grained geometric and relational fidelity in 3D assets. CLIP Score and Pick-a-Pic treat each rendered image independently and compare it to the same text prompt, without accounting for the viewpoint from which the image was rendered. As a result, assets with geometric inconsistencies—such as those suffering from the Janus problem—can receive artificially high scores, especially from favorable viewpoints. Although Text-to-3D Alignment is designed for 3D evaluation, it still struggles in edge cases where outputs are close to the prompt but miss key compositional or relational details. In contrast, the user study, which involves human assessment, is better equipped to capture these nuances, providing a more reliable measure of visual, structural, and semantic fidelity across complex scenes.
A.7 Additional Ablation Studies
@@
Azimuth Adjustment We ablate the impact of azimuth correction and object-view-dependent embeddings in Eq. (4). Without azimuth correction, objects exhibit noticeable disorientation. As illustrated in Figure 9, omitting object-view adjustments in the robot example leads to misplaced features—for instance, the robot’s eyes appear on the back of its head. This misalignment arises because the global front view of the entire composition does not align with the object’s canonical front. As a result, the losses $\mathcal{L}_{\text{obj}}$ and $\mathcal{L}_{\text{joint}}$ may optimize in conflicting directions, hindering convergence and structural coherence.
@@
Comparison with Stable Diffusion 2.1 We further analyze the impact of using SDS-based SD2.1 versus SDS-based SD3.5 in DecompDreamer. While SDS-based SD2.1 outperforms other SOTA methods in relationship modeling, it still struggles with sharpness and convergence speed. By replacing SDS with SDS, we significantly improve object detail and reduce optimization time, requiring only one-fourth of the iterations. As shown in Figure 15, SDS produces sharper edges and more refined object details compared to SDS.
@@
A.8 Sample usage of ChatGPT
A.9 More Comparisons
A.10 Failure Cases and Limitations
While DecompDreamer excels at generating 3D assets composed of distinct object types, it still exhibits several limitations. (1) As shown in Figure 14(a), for the prompt “a piglet playing a piano”, both the piglet and the piano suffer from the Janus problem. This arises due to the dominance of front-facing views in the training distribution of text-to-image models, limiting the model’s ability to generalize to uncommon or challenging viewpoints. To address this issue, incorporating high-fidelity control signals—such as reference 3D objects from datasets like Objaverse—can offer stronger guidance in handling such unseen views, thereby improving both realism and view consistency. (2) Figures 14(b) and (c) illustrate the model’s difficulty in handling complex actions (e.g., a bag with its zipper open) and decomposing a single entity into fine-grained subcomponents—such as segmenting a 3D human into semantically meaningful parts like the face, hands, or limbs. In such intricate compositional scenarios, textual prompts often lack the granularity required to effectively supervise the generation process and can introduce ambiguity when lifting 2D semantics into 3D space. These cases may benefit from higher-order supervision signals, such as reference images or spatial layouts.
Figure 14: Examples of failure cases. (a) DecompDreamer exhibits the Janus problem. (b) A case where DecompDreamer fails to model complex actions, such as a bag with its zipper open. (c) Difficulty in decomposing a single entity into fine-grained subcomponents—such as segmenting human body parts.
@@ A.10 Failure Cases and Limitations
Figure 16: More qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D Zhou et al. (2024), GraphDreamer Gao et al. (2024), Trellis Xiang et al. (2024), MVDream Shi et al. (2024), LucidDreamer Liang et al. (2024), Magic3D Lin et al. (2023) and HiFA Zhu et al. (2023).
Figure 17: More qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D Zhou et al. (2024), GraphDreamer Gao et al. (2024), Trellis Xiang et al. (2024), MVDream Shi et al. (2024), LucidDreamer Liang et al. (2024), Magic3D Lin et al. (2023) and HiFA Zhu et al. (2023).
Figure 18: More qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D Zhou et al. (2024), GraphDreamer Gao et al. (2024), Trellis Xiang et al. (2024), MVDream Shi et al. (2024), LucidDreamer Liang et al. (2024), Magic3D Lin et al. (2023) and HiFA Zhu et al. (2023).